it takes a boat 9 hours to go a distance of 270 miles against the current and 3 hours to make the return trip with the current. Find the rate of the current
Answer by stanbon(75887)

it takes a boat 9 hours to go a distance of 270 miles against the current and 3 hours to make the return trip with the current. Find the rate of the current
-------
Against current:
rate = 270/9 = 30 mph
-----
With current:
rate = 270/3 = 90 mph
---------------------------
Equations:
b + c = 90 mph
b - c = 30 mph
------------------
2b = 120
b = 60 mph
------------------
Solve for "c":
b + c = 90
60 + c = 90
current = 30 mph
=====================
Cheers,
Stan H.